# Bayesian Rank-Clustering

### v1

@@ 5.1 Sushi Preferences in Tohoku
We observe that policy option 2 (information campaigns) is strictly preferred to the rest among the population of survey respondents from Great Britain, whereas options 5 (funding research) and 3 (legal penalty) are rank-clustered last. The results indicates to policymakers that respondents in Great Britain strongly prioritize Option 2 in comparison to the rest, while pairs of Options 1 and 4 and Options 3 and 5, are, respectively, indistinguishable within each pair, with 1 and 4 being strongly preferred to 3 and 5. By rank-clustering similarly-preferred options, interpretation of the population’s preferences is simplified for policymakers.
5.4 2020-21 National Basketball Association Game Outcomes
@@
Finally, we analyze outcomes of 1,249 pairwise games in the 2020-21 season of the National Basketball Association (NBA). In this season, there were 30 teams, who each played every other team at least twice; 3 pairs of teams played 11 matches. We seek to estimate an overall ranking of teams that allows for potential equality in ranking.
@@
We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=0.1$ to encourage rank-clustering given the limited ordinal comparison data provided by pairwise matchups. Figure 9 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 9: Posterior rank-clustering probabilities (left) and posterior distributions of basketball teams (right) based on the results of all 2020-21 regular season games.
In this setting, the Rank-Clustered BTL model estimates an ordering of professional basketball teams with uncertain rank-clustering patterns. Note than in the left panel of Figure 9, the squares on the main diagonal are black (indicating probability 1) by default since each team is inherently rank-clustered with itself. An 80% posterior credible interval indicates that there are between 4 and 6 rank-clusters; every team has less than 0.015 probability of belonging to a singleton rank-cluster. However, rank-clusters in this setting are less certain than in previous examples. Most clearly, 5 teams (Detroit Pistons, Oklahoma City Thunder, Cleveland Cavaliers, Orlando Magic, and the Houston Rockets) appear to be rank-clustered for last place. Among these 5 teams, pairwise rank-clustering probabilities are at least 0.75. Instead of ranking these teams with granularity, we may simply infer by our model that they were each approximately the worst teams of the league in this season. Conversely at the top of the overall ranking, we observe that the top 5 teams (Utah Jazz, Phoenix Suns, Philadelphia 76ers, Milwaukee Bucks, and the Brooklyn Nets) all have a high posterior probability of rank-clustering. Although the Milwaukee Bucks were the ultimately this season’s league champions, our model suggests that these 5 teams were of roughly indistinguishable ability based on the game data. These summaries from our model, despite not accounting for the complexities of the sport, provide useful and interpretable summaries of the teams’ abilities across the season.
@@ 6 Discussion
In this paper, we proposed the Rank-Clustered Bradley-Terry-Luce model for estimating an overall ranking of objects with rank-clusters. The model employs the Bradley-Terry-Luce (BTL) family of distributions for ordinal comparison data and estimates model parameters in a Bayesian framework under the proposed Partition-based Spike-and-Slab (PSSF) prior. The model requires no pre-specification of the number or size of the rank-clusters (as in Piancastelli and Friel (2024)), nor specification of lasso-based penalty parameters (as in Masarotto and Varin (2012); Jeon and Choi (2018)). In a simulation study, we demonstrated the model’s ability to accurately and consistently estimate the relative worth of objects in a collection while simultaneously estimating rank-clusterings over objects. We used Rank-Clustered BTL on four real datasets under different types of ordinal comparison data.
@@
In contrast to the only other spike-and-slab based prior for parameter fusion Wu et al. (2021), PSSF prior we developed does not require a known parameter order. Visual inspection of the prior distribution makes obvious its connection to spike-and-slab: “spike” components correspond to parameter clusters and “slab” components correspond to independent parameters. Estimation of parameters under this model requires reversible jump MCMC. To overcome potentially slow or computationally-burdensome estimation in this setting, we proposed a computationally efficient Gibbs sampler that alternates between updating the partition of objects (based on the seminal work of Green (1995) and updating object-level worth parameters by adapting a data augmentation trick for standard Plackett-Luce models by Caron and Doucet (2012) and later adapted for Plackett-Luce mixtures by Mollica and Tardella (2017).
@@
A useful benefit of estimating parameter values and clusters in a single Bayesian framework is the avoidance of issues associated with selective inference (Taylor and Tibshirani, 2015) or more colloquially, double dipping (Kriegeskorte et al., 2009). Selective inference occurs when the same data is used twice in the process of model selection and/or estimation, e.g., to estimate some latent structure underlying the data and subsequently to estimate parameters conditional on that estimated structure. In our context, selective inference would occur if ordinal preference data was used first to identify rank-clusters and then used again to estimate worth parameter values conditional on those clusters. We note that selective inference occurs in the estimation of the related Clustered Mallows Model by Piancastelli and Friel (2024), which requires selecting the number and size of rank-clusters among objects before fitting the model. Selective inference often leads to invalid inference in part because uncertainty regarding the estimated clustering structure is not taken into account. However, Rank-Clustered Bradley-Terry-Luce models do not perform selective inference because parameter values and rank-clusters are estimated simultaneously. As such, our parameter estimates incorporate uncertainty across the posterior distributions of both the rank-clustering structure and the specific parameter values.
@@
Results from Rank-Clustered BTL models are useful in a variety of contexts. As noted in other fusion literatures on rankings, estimated overall rankings may be easier to understand and interpret when rank-clusters of objects are identified, as rank-clusters lead to fewer rank levels of objects to distinguish (Masarotto and Varin, 2012). In contexts where model results are used for prediction, such as in sports, estimating rank-clusters may improve predictive accuracy (Tutz and Schauberger, 2015). Similarly, estimating rank-clusters is important in the context of decision-making: In peer review, for example, rank-clusters can be beneficial for communicating uncertainty in the assessment of preferences and for better transparency in funding decisions. We might imagine a scenario where a government agency is only able to fund two grants, however, two grant proposals are rank-clustered in second place. In this case, rank-clustering can be used to communicate uncertainty in the relative quality of the top proposals. A potential danger is that under this uncertainty, decision makers may be tempted to resort to unfair tie-breaking methods, e.g., selecting the proposal with the most famous author. Instead, tie-breaking should occur based on a fairer or more principled method, such as a partial lottery (Fang and Casadevall, 2016; Roumbanis, 2019; Heyard et al., 2022).
@@
We list a few possible directions for future research. First, in this work we have not considered the level of interconnectedness among the assessed objects (e.g., if separate groups of judges assess completely distinct sets of objects). This is particularly relevant in the case of pairwise comparison data, in which some pairs of objects may never experience a head-to-head match-up. Second, the PSSF prior could be imposed as a prior for more complex BTL models or to other models entirely. In the former, the PSSF prior could be applied to preference learning via BTL distributions that incorporate covariates (e.g., Baldassarre et al. (2023); Gormley and Murphy (2010); Chapman and Staelin (1982)). In that case, the prior may be modified to permit covariate parameter estimation in addition to rank-clustering. In the latter case, the PSSF prior may be applied to regression for variable fusion, and its performance may be compared to other existing Bayesian variable fusion methods (e.g., Casella et al. (2010); Song and Cheng (2020); Shimamura et al. (2019)).
@@ Acknowledgements
Figure 19 displays stacked bar charts of the season record of each NBA team across the 2020-21 season.
Figure 19: Stacked bar charts of ranks received by each NBA team across the 2020-21 season. Winning = rank 1; losing = rank 2.
@@
Figure 20: Trace plot of K in the 2020-21 NBA season analysis
Figure 21: Trace plot of $\omega$ in the 2020-21 NBA season analysis

### v2

@@ 5.1 Sushi Preferences in Tohoku
Policy option 2 (information campaigns) is strictly preferred to the rest among the population of survey respondents from Great Britain, whereas options 5 (funding research) and 3 (legal penalty) are rank-clustered last. The results indicates to policymakers that respondents in Great Britain strongly prioritize Option 2 in comparison to the rest, while pairs of Options 1 and 4 and Options 3 and 5, are, respectively, indistinguishable within each pair, with 1 and 4 being strongly preferred to 3 and 5. By rank-clustering similarly-preferred options, interpretation of constituent preferences is simplified for policymakers.
5.4 2023-24 National Basketball Association Game Outcomes
@@
Last, we analyze outcomes of 1,230 games from the 2023-24 season of the National Basketball Association (NBA) of the United States of America (National Basketball Association, 2024). In this season, 30 teams each played 82 games, including between 2 and 5 games against every other team. We seek to estimate an overall ranking of teams that allows for potential equality in ranking.
@@
We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=1$ to encourage rank-clustering given the limited ordinal comparison data provided by pairwise matchups. Figure 9 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics are provided in the supplementary materials.
Figure 9: Primary results from Rank-Clustered BTL analysis of 2023-24 NBA data. Left: Posterior rank-clustering probabilities. Main diagonal displays posterior median estimate of worth parameter after normalization. Right: Posterior distributions of team-specific worth parameters.
In this setting, the Rank-Clustered BTL model estimates an ordering of professional basketball teams with uncertain rank-clustering patterns. Uncertain rank-clustering may result from two aspects of this application. First, pairwise comparisons provide little information in relation to partial or complete rankings, by construction. Second, game outcomes provide low signal measurements of team ability (Baumer et al., 2023). That is because many factors influence game outcomes, such as skill, home advantage, injuries, roster changes, and luck (Cai et al., 2019). Consistent with the low signal and limited information setting, an 80% posterior credible interval indicates that there are between 6 and 9 rank-clusters. Every team has less than 0.037 posterior probability of belonging to a singleton rank-cluster.
As seen in the left panel of Figure 9, 4 teams (Boston Celtics, Oklahoma City Thunder, Denver Nuggets, and Minnesota Timberwolves) appear to be rank-clustered for first place. Based on regular season data alone, our model suggests that these 4 teams were of roughly indistinguishable ability. Conversely, we observe that 6 teams (Toronto Raptors, San Antonio Spurs, Portland Trail Blazers, Charlotte Hornets, Washington Wizards, and Detroit Pistons) all have a high posterior probability of rank-clustering in last place. Instead of reporting the uncertain ranking of these teams with some granularity, we recommend to infer that these teams were the worst teams of the league in this season. These rank-clusters, despite not accounting for the complexities of the sport, provide useful and interpretable summaries of the teams’ abilities across the regular season. A similar analysis could be used in the future to predict postseason performance.
@@ 6 Discussion
In this paper, we proposed the Rank-Clustered Bradley-Terry-Luce model for estimating an overall ranking of objects with rank-clusters. The model employs the Bradley-Terry-Luce (BTL) family of distributions for ordinal comparisons. We proposed Partition-based Spike-and-Slab Fusion (PSSF) prior to estimates model parameters in a Bayesian framework. The model requires neither pre-specification of the number or size of the rank-clusters (improving upon Piancastelli and Friel (2024)), nor specification of lasso-based penalty parameters (improving upon Masarotto and Varin (2012); Jeon and Choi (2018); Hermes et al. (2024)). In a simulation study, we demonstrated the model’s ability to accurately and consistently estimate the relative worth of objects in a collection while simultaneously estimating rank-clusters. We used Rank-Clustered BTL on four real datasets under different types of ordinal comparison data.
@@
In contrast to the only other spike-and-slab based prior for parameter fusion Wu et al. (2021), PSSF prior we developed does not require a known parameter order. Visual inspection of the prior distribution makes obvious its connection to spike-and-slab: “spike” components correspond to parameter clusters and “slab” components correspond to independent parameters. Estimation of parameters under this model requires reversible jump MCMC. To overcome potentially slow or computationally-burdensome estimation in this setting, we proposed a computationally efficient Gibbs sampler. The sampler alternates between updating the partition of objects, based on the seminal work of Green (1995), and updating object-level worth parameters following a data augmentation trick for standard Plackett-Luce models by Caron and Doucet (2012) that was later adapted for Plackett-Luce mixtures by Mollica and Tardella (2017).
The proposed PSSF prior requires selecting hyperpriors for partitions, $f_{G}$, and the continuous values for each unique parameter, $f_{\nu}$. In this work, we specified $f_{G}\propto\text{Poisson}(K_{g}|\lambda)$ to be intentionally vague over the large space of partitions and $f_{\nu}\propto\text{Gamma}(a_{\gamma},b_{\gamma})$ based on conjugacy. However, alternative hyperpriors are available. A Negative Binomial or Beta Negative Binomial distribution for $f_{G}$ may be more appropriate when stronger prior knowledge of $K_{g}$ is available. If PSSF were to be applied to linear regression for parameter fusion, a Normal or $t-$distribution may be substituted for $f_{\nu}$.
@@
A useful benefit of estimating parameter values and clusters in a single Bayesian framework is the avoidance of issues associated with selective inference (Taylor and Tibshirani, 2015) or, more colloquially, double dipping (Kriegeskorte et al., 2009). Selective inference occurs when the same data is used twice in the process of model selection and/or estimation, e.g., to estimate some latent structure underlying the data and subsequently to estimate parameters conditional on that estimated structure. In our context, selective inference would occur if ordinal preference data was used first to identify rank-clusters and then used again to estimate worth parameter values conditional on those clusters. We note that selective inference occurs in the estimation of the related Clustered Mallows Model by Piancastelli and Friel (2024), which requires selecting the number and size of rank-clusters among objects before fitting the model. Selective inference often leads to invalid inference in part because uncertainty regarding the estimated clustering structure is not taken into account. However, Rank-Clustered Bradley-Terry-Luce models do not perform selective inference because parameter values and rank-clusters are estimated simultaneously. As such, our parameter estimates incorporate uncertainty across the posterior distributions of both the rank-clustering structure and the specific parameter values.
@@
Results from Rank-Clustered BTL models are useful in a variety of inferential contexts. As noted in other fusion literatures on rankings, estimated overall rankings may be easier to understand and interpret when rank-clusters of objects are identified, as rank-clusters lead to fewer rank levels of objects to distinguish (Masarotto and Varin, 2012). In contexts where model results are used for prediction, such as in sports, estimating rank-clusters may improve predictive accuracy (Tutz and Schauberger, 2015). Similarly, estimating rank-clusters is important in the context of decision-making: In peer review, for example, rank-clusters can be beneficial for communicating uncertainty in the assessment of preferences and for better transparency in funding decisions. We might imagine a scenario where a government agency is only able to fund two grants, however, two grant proposals are rank-clustered in second place. In this case, rank-clustering can be used to communicate uncertainty in the relative quality of the top proposals. A potential danger is that under this uncertainty, decision makers may be tempted to resort to unfair tie-breaking methods, e.g., selecting the proposal with the most famous author. Instead, tie-breaking should occur based on a fairer or more principled method, such as a partial lottery (Fang and Casadevall, 2016; Roumbanis, 2019; Heyard et al., 2022).
@@
We list a few possible directions for future research. First, in this work we have not considered the level of interconnectedness among the assessed objects (e.g., if separate groups of judges assess completely distinct sets of objects). This is particularly relevant in the case of pairwise comparison data, in which some pairs of objects may never experience a head-to-head match-up. Second, the PSSF prior could be imposed as a prior for more complex BTL models or to other models entirely. In the former, the PSSF prior could be applied to preference learning via BTL distributions that incorporate covariates (e.g., Baldassarre et al. (2023); Gormley and Murphy (2010); Chapman and Staelin (1982); Hermes et al. (2024)). In that case, the prior may be modified to permit covariate parameter estimation in addition to rank-clustering. In the latter case, the PSSF prior may be applied to regression for variable fusion, and its performance may be compared to other existing Bayesian variable fusion methods (e.g., Casella et al. (2010); Song and Cheng (2020); Shimamura et al. (2019)). Third, we notice that the PSSF prior bears some resemblance to a Dirichlet process prior (Escobar and West, 1995). The connection between Bayesian nonparametrics and Bayesian parameter fusion requires further study.
@@ Acknowledgements
Figure 19 displays stacked bar charts of the season record of each NBA team across the 2023-24 season.
Figure 19: Stacked bar charts of ranks received by each NBA team across the 2023-24 season. Winning = rank 1; losing = rank 2.
@@
Figure 20: Trace plot of K in the 2023-24 NBA season analysis
Figure 21: Trace plot of $\omega$ in the 2023-24 NBA season analysis